OpenStudy (anonymous):

If the common difference in an arithmetic sequence is 4, and the 20th term is 36, what is the first term? A. –40 B. 76 C. 40 D. 2

OpenStudy (anonymous):

The answer is -40. I will tell you why.

OpenStudy (anonymous):

Let a[1]=x be the first term, then a[2]=4 + x a[3]=8+x; a[4]= 12 +x a[20]=76 +x=36 x=-40

OpenStudy (anonymous):

a[n]= 4(n-1) +a[1] a[20]= 4(19) +x= 76 +x